# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

### v1

@@ 3. Research methodology
We have conducted an SLR on WaaS broker according to the guidelines of (Petersen et al., 2015; Kitchenham et al., 2009). The used SLR methodology has four phases (Khorasani et al., 2024), which are elaborated in the following.
3.1. Specifying the scope and research questions
@@
Research questions (RQs) are the foundation of a systematic study. In fact, an SLR is an effort to answer RQs. In this study, we have defined 5 research questions as follows:
@@
Our review covers a period of 20 years, from 2005, the advent of cloud computing, to the end of 2024. We have considered research works that explicitly addressed the design of a WaaS platform for cloud computing environments, or proposed a multiple workflow scheduling algorithm from the client-side perspective in cloud environments.
3.2. Planning the search process
@@
These search strings were applied to the Title, Abstract, and Keywords sections of the papers to retrieve all relevant papers. All extracted papers and their venues (i.e., journals or conferences) were evaluated against predefined criteria (see the next sub-section), and the papers meeting those criteria were added to the set of included papers, which was then sent to the snowballing step.
@@
3.3. Planning the study selection process
The first step of this phase is determining the criteria for venue inclusion/exclusion to ensure that high-quality venues are selected. For journals, those indexed in the Journal Citation Reports (JCR) (ranks Q1-Q4) are included. For conferences, those indexed in the ICORE conference ranking (Computing Research and Education Association of Australasia, [n. d.]) (ranks A, B, and C) are added to the list of included conferences. All extracted studies that are not published in the included venues are discarded from further consideration.
The next step is defining the study selection strategy. Initially, the relevance of a paper is evaluated by reviewing its title, abstract, and keywords. If these elements do not provide sufficient information to make a definitive decision about its relevance, the full text of the paper is reviewed. For a paper to be included, it should explicitly design a workflow as a service or propose a multiple workflow scheduling algorithm for the cloud environment. Papers proposing scheduling algorithms for a single workflow, or those designed for other distributed environments such as clusters, grids, edge, or fog, have been excluded. Furthermore, papers that propose scheduling algorithms from the cloud provider’s point of view, which schedule on physical machines (such as (Wang et al., 2023, 2024; Fan et al., 2023a)), are excluded. A few papers have also been excluded because, although they claimed to be designed for the cloud environment, they did not consider important properties of the cloud, such as elasticity and scalability.
3.4. Conducting the SLR and analyzing results
After establishing the strategies, we proceeded with the SLR by implementing these strategies. As a result, we identified 43 venues (31 journals and 12 conferences), along with 74 studies that were included, which are detailed in the provided Excel supplementary file. Figure 2 presents a brief overview of the studies included per year and per publisher. As seen in Figure 2(a), the number of published papers shows an increasing trend in recent years, reaching its peak in 2024. This suggests that researchers have paid more attention to this topic in recent years. Please note that the figure starts from 2014, as only two papers were published before in 2009 and 2011. Figure 2(b) illustrates that IEEE journals and conferences have the highest number of papers in this area, followed by Springer and Elsevier. The top three venues with the most published papers are The Journal of Supercomputing with 7, and Future Generation Computer Systems and IEEE Transactions on Services Computing both with 6 published papers.
@@ 3.4. Conducting the SLR and analyzing results
Figure 2. Number of included papers
The final step involves analyzing and classifying the data to address the research questions of the SLR. The primary objectives of this SLR are to identify the taxonomy of the WaaS broker and to classify the included papers within this taxonomy. After a thorough review of the included papers, and based on the proposed architecture of the WaaS broker in Figure 1, a taxonomy was developed for each of the four components of a WaaS broker. All included papers were carefully analyzed by two authors and classified according to the taxonomy. Any disagreements were reviewed and resolved by all authors. The results of the analysis are presented in Section 4.
4. Taxonomy and survey
In this section, we elaborate on the proposed taxonomy and then survey the related work based on this taxonomy. The taxonomy has been designed around the different components of a WaaS broker, and we will discuss it in the following four sub-sections, each focusing on a specific component. To conserve space in this SLR, a list of all included articles along with their classification results has been provided as a supplementary file in the form of an Excel document. This also facilitates easier searching and review of the included articles for readers.
4.1. Negotiation and pricing taxonomy
@@
[…]
Figure 3. Negotiation and pricing taxonomy
@@
4.1.1. QoS Requirements
Users have varying requirements regarding quality of service criteria (e.g., execution time, cost, privacy, reliability, etc.; see Section 4.2.5). These requirements can be classified into two categories. The first category is Multi-criteria, where the user seeks to optimize multiple criteria simultaneously. In this case, the broker faces a multi-criteria or multi-objective optimization problem, which can be addressed by finding Pareto Optimal solutions or by optimizing a weighted sum of the criteria. Many studies in this category focus on minimizing both the makespan and cost of workflow execution (Wang et al., 2019; Xia et al., 2023; Li et al., 2022a; Li et al., 2024d), while others consider additional objectives, such as makespan, cost, and energy at the same time (Tarafdar et al., 2023). For example, Sun et al. (Sun et al., 2023) addressed a tri-objective multi-workflow scheduling problem in a hybrid cloud, aiming to minimize the total tardiness (i.e., the maximum between zero and the makespan of the workflow minus its deadline), total energy consumption in the private cloud, and the total renting cost in the public cloud. Similarly, Li et al. (Li et al., 2024b) considered the optimization of three objectives simultaneously: makespan, cost, and fairness.
On the other hand, the Constrained class focuses on optimizing one criterion while treating the other criteria as constraints. Two common constraints in this class are deadline-constrained and budget-constrained requirements. In a deadline-constrained scenario, the broker aims to meet the user-defined deadline while optimizing another criterion, typically minimizing the cost of workflow execution (Taheri et al., 2023; Rodriguez and Buyya, 2018). In a budget-constrained scenario, the broker seeks to keep the execution cost within the user-defined budget while optimizing other criteria, usually minimizing the workflow execution time (Li et al., 2023b, 2024a). Some studies address both deadline and budget constraints simultaneously (Arabnejad and Barbosa, 2017a, b; Rezaeian et al., 2019). Finally, there are also studies that optimize multiple objectives while ensuring that certain constraints are met. For example, Karmakar et al. (Karmakar et al., 2024) optimized resource utilization and rental costs while ensuring that the workflows met their deadlines.
@@ 4.1.1. QoS Requirements
Figure 4. Number of articles on QoS Requirements and Workflow Arrival.
4.1.2. Workflow Arrival
@@
Workflow arrival patterns can be categorized into four classes. In the online (or random) arrival pattern, workflows arrive at random times, with no predefined schedule (Chen et al., 2021; Liu et al., 2021; Liu et al., 2024; Taghinezhad-Niar et al., 2022). As a result, the broker must be prepared to handle workflow arrivals at any time. In contrast, batch arrival occurs when a set of workflows arrives together and is sent to the scheduler to be executed simultaneously (Wu and Yuandou, 2018; Zhou et al., 2024; Li et al., 2024c). Some articles that claim to adopt an online arrival pattern actually aggregate incoming workflows within a time interval and send them to the scheduler as a batch. These articles are classified under the batch category in this review (Li et al., 2022d; Li et al., 2024d).
@@
The third category is ensemble arrival, which is similar to the batch arrival pattern but with a key distinction: all workflows in an ensemble share the same structure but have different sizes and input data (Zolfaghari and Abrishami, 2022; Malawski et al., 2015; Chen et al., 2024b; Jiang et al., 2015). Typically, users specify QoS requirements, such as deadlines and budgets, for the entire ensemble rather than for individual workflows. The final category is periodic arrival, where multiple workflows are released at regular intervals, with each having a defined period between consecutive instances. Generally, an instance of a workflow must complete before the next one is released, with the release of the next instance acting as a deadline for the previous one. Only two of the reviewed articles address this arrival pattern (Taheri et al., 2023; Chen et al., 2020a), and both consider using the reserved pricing model for running periodic workflows.
4.1.3. Negotiation
@@
Negotiation is the process where the user and the WaaS broker agree on an SLA, such as pricing or execution time. For instance, in a deadline-constrained case, the user may specify a desired deadline for workflow completion, and the broker and user can then agree on a price. Negotiation can be divided into two types: static and dynamic. In static negotiation, the user specifies their QoS requirements, and the broker checks if it can fulfill them. If it can not, the request is declined, but if it can, the broker sends back a price for the user to accept or reject. However, standards like WS-Agreement Negotiation (Battré et al., 2010) enable the broker to negotiate QoS and price dynamically with users. For example, in a deadline-constrained scenario, if the broker cannot meet the specified deadline, it might propose a feasible deadline to the user. Or if the price is too high, the broker could offer a lower price in exchange for a more flexible deadline. Unfortunately, dynamic negotiation has not been employed by any of the reviewed articles, and only one article has implemented static negotiation. Thus, developing negotiation strategies could be a promising area for future research in WaaS broker design.
Among the reviewed studies, only Taghavi et al. (Taghavi et al., 2023) proposed a static negotiation module for a WaaS broker. They designed a deadline-constrained broker where the user submits a workflow with a specified deadline. The broker calculates the workflow’s critical path and accepts it if the deadline is sufficiently distant from the critical path (based on parameters set by the broker’s administrator). The broker then calculates the price for execution, considering factors like the deadline, resource availability, costs, and the broker’s profit margin. This price is presented to the user for approval or rejection. The negotiation component shows significant potential as a subject for future research directions.
4.1.4. Pricing
@@
Pricing is the process of determining the cost of the services offered by the WaaS broker. To set the price, the broker must account for the actual cost of executing the workflow, its overhead costs (such as penalties for violating the SLA), its profit margin, and the user’s benefits. Typically, the price should be lower than the cost of the user executing the workflow directly by leasing resources from the cloud provider. When using on-demand instances, the broker is expected to reduce costs by reusing leased resources for multiple workflows. However, the broker can significantly lower execution costs by using reserved and spot instances. In this case, the pricing is determined based on broker-side factors, such as available idle reserved instances or the current spot instance price, along with client-side factors, such as the specified deadline for the workflow.
@@
Pricing can be approached in two ways: pre-pricing and post-pricing. In pre-pricing, the broker sets the price for executing the workflow before it begins. This involves predicting the execution cost by examining the workflow’s structure, its required QoS, and the current availability and pricing of resources. Specifically, when using reserved or spot instances, the broker must assess the idle capacity of reserved instances and estimate the future cost of spot instances used for execution. The broker then uses this predicted cost to establish a final price, factoring in hidden costs and its profit. Pre-pricing is also employed during the negotiation phase to finalize the price with the user ahead of execution. In contrast, post-pricing involves determining the final price based on the actual cost incurred during workflow execution. While post-pricing offers greater accuracy and is more beneficial for the broker, users typically prefer knowing the price in advance to decide whether to accept or reject the offer.
@@
Although the majority of the papers reviewed focus on the execution cost of workflows, either by minimizing the cost or staying within the user-defined budget, only one paper addresses the pricing policies of WaaS brokers. Taghavi et al. (Taghavi et al., 2023) introduce a post-pricing policy, called constant profit, and two pre-pricing policies: constant discount and prediction-based, for a deadline-constrained WaaS broker that utilizes a combination of on-demand and spot instances for workflow execution. In the constant profit model, after completing the workflow, the broker calculates the total execution cost and adds a fixed profit margin. In the constant discount approach, the broker estimates the workflow’s execution cost using on-demand instances and then applies a discount based on past experience with spot instances before presenting the final price to the user. The prediction-based pricing method involves the broker estimating the cost of execution ahead of time, using the current prices of on-demand instances and the anticipated cost of spot instances during the workflow’s runtime. These pricing strategies highlight a promising area for further development in WaaS brokers.
4.2. Task scheduling taxonomy
@@
[…]
Figure 5. Task scheduler taxonomy.
4.2.1. Task Type
@@
The task type defines the degree of flexibility in the task’s execution time. Some studies consider tasks to be rigid, meaning they require a fixed amount of computational resources, with execution time remaining (nearly) constant on these resources (Taghavi et al., 2023; Wang et al., 2014). However, most studies focus on flexible tasks, which exhibit varying performance depending on the resources used. To represent flexibility, some studies assume the execution time of each task on each resource is pre-determined through techniques like profiling (Wang et al., 2019). Others assume that each task has a known execution time on the fastest resource, with other resources assigned a weight reflecting their relative computational power. Based on this weight, the task’s execution time on any resource can be derived (Taghinezhad-Niar and Taheri, 2023; Liu et al., 2024). Lastly, some studies estimate the task size in Million Instructions (MI) and the resource’s CPU capacity in MIPS (Million Instructions Per Second), calculating the task’s execution time by dividing these two values (Rodriguez and Buyya, 2018; Tarafdar et al., 2023).
4.2.2. Execution Time Estimation
@@
Most proposed scheduling algorithms rely on an estimation of each task’s execution time. Many studies assume a fixed execution time (Arabnejad and Barbosa, 2017b; Esteves and Veiga, 2015), which is impractical in dynamic environments like the cloud, where resource performance can be unpredictable. To address this, some researchers adopt probabilistic execution times for workflow tasks, modeling task execution time as a stochastic variable with an independent normal distribution characterized by a specific mean and variance (Ma et al., 2021; Liu et al., 2021; Xu et al., 2023). For instance, Chet et al. (Chen et al., 2021) model both task execution time and data transfer time among tasks as random variables and use their quantiles for approximation. A few studies manage execution time uncertainty using fuzzy logic. Zhu et al. (Zhu et al., 2020) represent task execution times and data transfer times using triangular fuzzy numbers. Based on these representations, they calculate fuzzy start and completion times for each task. They then introduce a list-based heuristic algorithm, Fuzzy Task Scheduling (FTS), which uses fuzzy priorities to schedule tasks effectively.
4.2.3. Task Concurrency

### v2

@@ 3 Research Methodology
We have conducted an SMS on WaaS broker according to the guidelines of [127, 92]. The used SMS methodology has four phases [90], which are elaborated in the following.
3.1 Specifying the Scope and Research Questions
@@
Research questions (RQs) are the foundation of a systematic study. In fact, an SMS is an effort to answer RQs. In this study, we have defined 5 research questions as follows:
@@
Our review spans 21 years, from the advent of cloud computing in 2005 to the first half of 2025. It considers studies that explicitly design a WaaS platform for cloud environments or propose algorithms for the MWSP from the client-side perspective.
3.2 Planning the Search Process
@@
These search strings were applied to the Title, Abstract, and Keywords sections of the papers to retrieve all relevant papers. All extracted papers and their venues were evaluated against predefined criteria, and the papers meeting those criteria were added to the set of included papers, which was then sent to the snowballing step.
@@
3.3 Planning the Study Selection Process
The first step of this phase is determining the criteria for venue inclusion/exclusion to ensure that high-quality venues are selected. For journals, those indexed in the Journal Citation Reports (JCR) are included. For conferences, those indexed in the ICORE conference ranking [43] are added to the list of included conferences. All extracted studies that are not published in the included venues are discarded from further consideration.
The next step is to define the study selection strategy. Relevance was first assessed by screening each paper’s title, abstract, and keywords. When these elements did not provide sufficient information, the full text was reviewed. To be included, a study had to explicitly design a workflow as a service or propose a multiple workflow scheduling algorithm for the cloud environment. Studies focusing on single-workflow scheduling or targeting other distributed environments—such as clusters, grids, edge, or fog—were excluded (e.g., [40], which addresses a hybrid cloud–edge architecture). We also restricted the scope to articles that approach the MWSP from the client-side (or broker-side), excluding those that address it from the provider-side, such as scheduling tasks on physical machines (e.g., [175, 173, 55, 174]). Finally, studies limited to specific workflow types, such as simple pipelines rather than general workflows modeled as DAGs, were excluded (e.g., [155, 150]).
3.4 Conducting the SMS and Analyzing Results
@@
Figure 2: Number of included papers
We conducted the SMS following established strategies, identifying 49 venues (34 journals and 15 conferences) and 87 studies, all detailed in Appendix B. Figure 2 presents a brief overview of the studies included per year and per publisher. As shown in Figure 2(a), the number of published papers has followed an upward trend in recent years, peaking in 2024 (with 2025 covering only the first half). The figure starts from 2014, as only two papers were published earlier (2009 and 2011). Figure 2(b) shows IEEE leads in publications, followed by Springer and Elsevier. The top three venues are The Journal of Supercomputing with 8, Future Generation Computer Systems with 7, and IEEE Transactions on Services Computing with 6 published papers.
The final step involves analyzing and classifying the data to address the research questions of the SMS. The primary objectives of this SMS are to identify the taxonomy of the WaaS broker and to classify the included papers within this taxonomy. After a thorough review of the included papers, and based on the proposed architecture of the WaaS broker in Figure 1, a taxonomy was developed for each of the four components of a WaaS broker. All included papers were carefully analyzed by two authors and classified according to the taxonomy. Any disagreements were reviewed and resolved by all authors. The results of the analysis are presented in Section 4.
4 Taxonomy and survey
In this section, we elaborate on the proposed taxonomy and then survey the related work accordingly. The taxonomy is structured around the key components of a WaaS broker, which are discussed in the following four sub-sections, each dedicated to a specific component. To conserve space in this SMS, all included articles, along with their classifications under the proposed taxonomy, are provided in Appendix B as an Excel file. This format enables readers to conveniently search the articles and generate customized statistics and charts. In addition, Appendix A reviews real-world adoption and industrial implementations of WaaS brokers.
4.1 Negotiation and Pricing Taxonomy
@@
Figure 3: Negotiation and pricing taxonomy
@@
4.1.1 QoS Requirements
Users have diverse requirements concerning QoS criteria, such as execution time, cost, privacy, reliability, and others (see Section 4.2.5). These requirements can generally be divided into two categories. The first category is multi-criteria, where the user expects multiple objectives to be optimized simultaneously. In this case, the broker is faced with a multi-objective optimization problem, which can be solved either by identifying Pareto-optimal solutions or by optimizing a weighted sum of the objectives. A large portion of studies in this category have concentrated on minimizing both the makespan and the execution cost of workflows [176, 185, 100, 101]. Other studies have extended the optimization to include more objectives simultaneously, e.g., makespan, cost, and energy consumption [165], or makespan, cost, and fairness [97].
On the other hand, the constrained class focuses on optimizing one criterion while treating the other criteria as constraints. Two common constraints in this class are deadline-constrained and budget-constrained requirements. In a deadline-constrained scenario, the broker aims to meet the user-defined deadline while optimizing another criterion, typically minimizing the cost of workflow execution [163, 135]. In a budget-constrained scenario, the broker seeks to keep the execution cost within the user-defined budget while optimizing other criteria, usually minimizing the workflow execution time [103, 99]. Some studies address both deadline and budget constraints simultaneously [15, 16, 131]. Finally, there are also studies that optimize multiple objectives while ensuring that certain constraints are met. For example, Karmakar et al. [87] optimized resource utilization and rental costs while ensuring that the workflows met their deadlines. Figure 4(a) indicates that the majority of previous studies adopted constrained QoS requirements.
@@ 4.1.1 QoS Requirements
Figure 4: Number of articles on QoS Requirements and Workflow Arrival.
4.1.2 Workflow Arrival
@@
Workflow arrival patterns can be categorized into four classes. In the online (or random) arrival pattern, workflows arrive at random times, with no predefined schedule [35, 110, 111, 160]. As a result, the broker must be prepared to handle workflow arrivals at any time. In contrast, batch arrival occurs when a set of workflows arrives together and is sent to the scheduler to be executed simultaneously [184, 198, 98]. Some articles that claim to adopt an online arrival pattern actually aggregate incoming workflows within a time interval and submit them to the scheduler as a batch [104, 101]. In this review, such approaches are classified under the batch category.
@@
The third category is ensemble arrival, which is similar to the batch arrival pattern but with a key distinction: all workflows in an ensemble share the same structure but have different sizes and input data [202, 114, 38, 82, 147]. Users usually specify QoS requirements, such as dedaline and budget, for the entire ensemble, not individual workflows, making scheduling and provisioning more challenging than in batch arrivals. The final category is periodic arrival, where multiple workflows are released at regular intervals, with each having a defined period between consecutive instances. Generally, an instance of a workflow must complete before the next one is released, with the release of the next instance acting as a deadline for the previous one. Only two of the reviewed articles address this arrival pattern [163, 37], and both consider using the reserved pricing model for running periodic workflows. Figure 4(b) demonstrates that the online pattern appears most frequently in prior research, with the batch arrival pattern being the second most common.
4.1.3 Negotiation
@@
Negotiation is the process where the user and the WaaS broker agree on an SLA, such as pricing or execution time. For instance, in a deadline-constrained case, the user may specify a desired deadline for workflow completion, and the broker and user can then agree on a price. Negotiation can be divided into two types: static and dynamic. In static negotiation, the user specifies their QoS requirements, and the broker checks if it can fulfill them. If it can not, the request is declined, but if it can, the broker sends back a price for the user to accept or reject. However, standards like WS-Agreement Negotiation [22] enable the broker to negotiate QoS and price dynamically with users. For example, in a deadline-constrained scenario, if the broker cannot meet the specified deadline, it might propose a feasible deadline to the user, at a lower price.
Dynamic negotiation has not been explored in any of the reviewed works; only Taghavi et al. [159] proposed a static negotiation module for a WaaS broker. Their approach considers deadline-constrained workflows: the user submits a workflow with a deadline, and the broker evaluates its feasibility by calculating the critical path. If the deadline is achievable, the broker estimates the execution price based on factors such as the deadline, resource availability, costs, and the broker’s profit margin, which is then offered to the user for approval or rejection. While this represents an initial step, negotiation—particularly dynamic negotiation strategies—remains an open and promising research direction (see Section 5.2).
4.1.4 Pricing
@@
Pricing is the process of determining the cost of the services offered by the WaaS broker. To set the price, the broker must account for the actual cost of executing the workflow, its overhead costs (such as penalties for violating the SLA), its profit margin, and the user’s benefits. Typically, the price should be lower than the cost of the user executing the workflow directly by leasing resources from the CSP. This aim can be achieved through efficient reuse of leased resources across multiple workflows, combined with cost-aware provisioning strategies that leverage lower-cost reserved and spot instances.
@@
Pricing can be approached in two ways: pre-pricing and post-pricing. In pre-pricing, the broker sets the price for executing the workflow before it begins. This involves predicting the execution cost by examining the workflow’s structure, its required QoS (e.g., deadline), and the current availability and costs of resources. Specifically, when using reserved or spot instances, the broker must assess the idle capacity of reserved instances and estimate the future cost of spot instances used for execution. The broker then uses this predicted cost to establish a final price, factoring in hidden costs and its profit. Pre-pricing is also employed during the negotiation phase to finalize the price with the user ahead of execution. In contrast, post-pricing involves determining the final price based on the actual cost incurred during workflow execution. While post-pricing offers greater accuracy and is more beneficial for the broker, users typically prefer knowing the price in advance to decide whether to accept or reject the offer.
@@
Although the majority of the papers reviewed focus on the execution cost of workflows, either by minimizing the cost or staying within the user-defined budget, only one paper addresses the pricing policies of WaaS brokers. Taghavi et al. [159] introduce a post-pricing policy, called constant profit, and two pre-pricing policies: constant discount and prediction-based, for a deadline-constrained WaaS broker that utilizes a combination of on-demand and spot instances for workflow execution. In the constant profit model, after completing the workflow, the broker calculates the total execution cost and adds a fixed profit margin. In the constant discount approach, the broker estimates the workflow’s execution cost using on-demand instances and then applies a discount based on past experience with spot instances before presenting the final price to the user. The prediction-based pricing method involves the broker estimating the cost of execution ahead of time, using the current prices of on-demand instances and the anticipated cost of spot instances during the workflow’s runtime. These pricing strategies highlight a promising area for further development in WaaS brokers.
4.2 Task Scheduling Taxonomy
@@
Figure 5: Task scheduler taxonomy.
4.2.1 Task Type
@@
The task type defines the degree of flexibility in the task’s execution time. Some studies consider tasks to be rigid, meaning they require a fixed amount of computational resources, with execution time remaining (nearly) constant on these resources [159, 172]. However, most studies focus on flexible tasks, which exhibit varying performance depending on the resources used. To represent flexibility, some studies assume the execution time of each task on each resource is pre-determined through techniques like profiling [176]. Others assume that each task has a known execution time on the fastest resource, with other resources assigned a weight reflecting their relative computational power. Based on this weight, the task’s execution time on any resource can be derived [161, 111]. Lastly, some studies estimate the task size in Million Instructions (MI) and the resource’s CPU capacity in MIPS (Million Instructions Per Second), calculating the task’s execution time by dividing these two values [135, 165].
4.2.2 Execution Time Estimation
@@
Most proposed scheduling algorithms rely on an estimation of each task’s execution time. Many studies assume a fixed execution time [16, 53], which is impractical in dynamic environments like the cloud, where resource performance can be unpredictable. To address this, some researchers adopt probabilistic execution times for workflow tasks, modeling task execution time as a stochastic variable with an independent normal distribution characterized by a specific mean and variance [112, 110, 186]. For instance, Chet et al. [35] model both task execution time and data transfer time among tasks as random variables and use their quantiles for approximation. A few studies manage execution time uncertainty using fuzzy logic. Zhu et al. [201] represent task execution times and data transfer times using triangular fuzzy numbers. Based on these representations, they calculate fuzzy start and completion times for each task. They then introduce a list-based heuristic algorithm, Fuzzy Task Scheduling (FTS), which uses fuzzy priorities to schedule tasks effectively.
4.2.3 Task Concurrency